OpenStudy (anonymous):

I NEED HELP! PLEASE!! this two are the LAST one!! factorize the following and solve for X: a) x(x-1)+3x=0 b)-x^3-5x^2+36x=0

OpenStudy (radar):

\[\]\[x ^{2}-x+3x=0\]

OpenStudy (anonymous):

Answer it quickly PLEASE! i have 5 mins o my hand now

OpenStudy (radar):

\[x ^{2}+2x=0\] \[x(x+2) =0\] x=0 x=-2

OpenStudy (anonymous):

which one is this?

OpenStudy (anonymous):

OH THANKS!! the next one? please? quick!!

OpenStudy (radar):

b.)\[-x ^{3}-5x ^{2}+36x=0\] \[-x(x ^{2}+5x-36)=0\] \[-x(x+9)(x-4)=0\] x=0 x=-9 x=4

OpenStudy (radar):

What say? Raka

OpenStudy (anonymous):

You are my lucky charm DUDE! Thank You! :D

OpenStudy (radar):

ur welcome